OpenStudy (anonymous):

integrate ln(4+x^2)dx

OpenStudy (blurbendy):

x(ln|x^2+4|-2) + 4arctan(x/2) + C

OpenStudy (anonymous):

using integration by parts

OpenStudy (anonymous):

let u = ln(4 +x^2) --->du = 2x/4+x^2 dv = dx ---> v = x. Replace everything into the formula you can have 2 parts. I just talk about the second part. which is \[\int\limits \frac{ 2x^2 }{ 4+x^2 }dx\]

OpenStudy (anonymous):

make a long division to get \[\int\limits2 +\frac{ 8 }{ x^2+4 }\]

OpenStudy (anonymous):

then, everything is easy. Hope this help

OpenStudy (anonymous):

yeah i was just stuck on the long division part thanks!

OpenStudy (anonymous):

Hi friend , I am sorry for the last part, It is negative , not positive, I mean -8/x^2+4. sorry for the mistake

OpenStudy (anonymous):

oh okay thank you

OpenStudy (anonymous):

you're welcome